OpenStudy (mimijola):

Find the distance between the points to the nearest tenth A(-1,5), B(0,4)

pooja195 (pooja195):

Hi there! Welcome to OpenStudy! :) Do you have any ideas as to where to start?

OpenStudy (mimijola):

Hi! Thank you! Well, based on my book I see that I have to use the distance formula but when I did the answer didn't come out right :(

pooja195 (pooja195):

Hmm.. Can you show me the work you did so far?

OpenStudy (mimijola):

Ok so I put (-1,5) (x1, y2) (0,4) (x2, y2) d=sqrt (0-(-1))^2 + (4-5)^2

OpenStudy (marziman):

(x^2-x^1) + (y^2-y^1)

OpenStudy (marziman):

Plug your values into that formula.

OpenStudy (mimijola):

=sqrt(1)^2 + (-1)^2 =sqrt 1+(-1) =0 but thats not right :(

OpenStudy (marziman):

Find the squareroot of 20 for your answer.

pooja195 (pooja195):

Ok let's work this out again, Here's the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We plug in our values \[\huge d=\sqrt{(0-(-1))^2+(4-5)^2}\] Continue from here

OpenStudy (marziman):

It's quite simple when you use the formula since the formula (x^2-x^1) + (y^2-y^1) is very straight forward. (-1^2-0^1)+(5^2-4^1) = 20

OpenStudy (marziman):

From there on you'd find the square root

OpenStudy (mimijola):

=sqrt(1)^2 + (-1)^2?

OpenStudy (marziman):

You don't have to solve the formula under the SQRT sign first @Mimijola, you can just solve the equation first then take the solution and get the sqrt

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Mimijola =sqrt(1)^2 + (-1)^2? \(\color{#0cbb34}{\text{End of Quote}}\) correct so far continue :)

OpenStudy (marziman):

Mimi I led you to the answer, what is the sqrt of 20?

OpenStudy (mimijola):

4.4.....

OpenStudy (marziman):

Yes

OpenStudy (marziman):

4.472136 to be specific.

OpenStudy (mimijola):

So that is the answer?

pooja195 (pooja195):

That's incorrect.

OpenStudy (mimijola):

My book says 1.4

pooja195 (pooja195):

That's the right answer but do you know how to get it?

OpenStudy (mimijola):

No, thats what Im confused about.

pooja195 (pooja195):

Ok let's work this out again, Here's the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] We plug in our values \[\huge d=\sqrt{(0-(-1))^2+(4-5)^2}\] Continue from here Bare in mind that 2 minus signs next to each other make a positive \[\huge d=\sqrt{(0+1)^2+(4-5)^2}\] Continue it

OpenStudy (marziman):

Whoa whoa, -1 should come before 0 since -1 is X1 being that it's A

pooja195 (pooja195):

Consider rereading the formula @marziman

OpenStudy (mimijola):

d=sqrt 1^2 + (-1)^2?

pooja195 (pooja195):

That good :) So we have \[\huge d=\sqrt{(1)^2+(-1)^2}\]

pooja195 (pooja195):

Keep going!

OpenStudy (mimijola):

\[\sqrt{1+(-1)}\]

OpenStudy (mimijola):

?

pooja195 (pooja195):

Not quite... -1^2=-1 *-1 =?

OpenStudy (mimijola):

0? or 1?

pooja195 (pooja195):

Well we are multiplying so which is it?

OpenStudy (mimijola):

1

pooja195 (pooja195):

Good :) so we have \[\huge~\rm~\sqrt{1+1}\] 1+1=?

OpenStudy (mimijola):

2

OpenStudy (mimijola):

Right?

pooja195 (pooja195):

Yep \[\huge~\rm~\bf~\sqrt{2}=? \]

OpenStudy (mimijola):

Yayyyy!! 1.4....

OpenStudy (mimijola):

Thank you!!!

OpenStudy (mimijola):

May I have your help with another one?

pooja195 (pooja195):

You're welcome :) And sure we can do another one

OpenStudy (mimijola):

Ok, thanks! Question: AB has endpoints A(-3,2) and B(3,-2) Find the coordinates of the midpoint AB

pooja195 (pooja195):

Here's the midpoint formula \[\huge~\rm~\frac{ x_2+x_1 }{ 2},\frac{ y_2+y_1 }{ 2 } \] We plug in our values \[\huge~\rm~\frac{ 3+(-3) }{ 2},\frac{ -2+2 }{ 2 } \]

OpenStudy (mimijola):

0/2, 0/2?

pooja195 (pooja195):

Well almost what is 0/2

OpenStudy (mimijola):

0?

pooja195 (pooja195):

Yep so the coordinates are?

OpenStudy (mimijola):

0,0?

pooja195 (pooja195):

Yep :)

OpenStudy (mimijola):

Omg Thank you very much! You saved my life :0

pooja195 (pooja195):

You're welcome! :)